OpenStudy (inkyvoyd):

Is it possible to solve the differential equation mr''(t)=-GMmr^-2 for r(t), t?

OpenStudy (loser66):

@wio

OpenStudy (inkyvoyd):

I can't seem to even get a numeric solution on wolfram alpha or mathematica because I just can't use the syntax correctly

OpenStudy (anonymous):

Well, I know the equation exists and it is a real valued function.

OpenStudy (anonymous):

function^

OpenStudy (loser66):

I think of reduction of order.

OpenStudy (anonymous):

But I don't know it is an elementary function.

OpenStudy (anonymous):

It looks separable.

OpenStudy (loser66):

to reduce the order from 2 to 1 by letting u = r' t

OpenStudy (anonymous):

But then again, it isn't first order.

OpenStudy (loser66):

sure, that's why I use that method

OpenStudy (anonymous):

\[ m\frac{d^2r}{(dt)^2}=G\frac{Mm}{r^2} \\ r^2\frac{d^2r}{(dt)^2}=GM \]

OpenStudy (loser66):

minus sign from the RHS, but then?

OpenStudy (anonymous):

What happens if you do it your way?

OpenStudy (inkyvoyd):

http://www.wolframalpha.com/input/?i=y%28x%29^2*y%27%27%28x%29%3Dc for some reason wolfram alpha won't give me an answer if I put a negative sign...

OpenStudy (loser66):

I didn't do, I tagged you , right? because I am not sure about it. ok, let me try

OpenStudy (anonymous):

Maybe intergration by parts @Loser66

OpenStudy (loser66):

to me, I let u = r'(t) u' = r"(t) = \(\dfrac{-GM}{r^2}\) then du =\(\dfrac{-GM}{r^2}\)dr integral both sides to get u

OpenStudy (inkyvoyd):

oh, to make life harder, I'm given r(0)=149,600,000,000, r'(0)=0, G=grav constant, M equals mass of the sun

OpenStudy (loser66):

@inkyvoyd numbers doesn't matter, plug them at the last step

OpenStudy (inkyvoyd):

@Loser66 if numbers don't matter then why is it that wolfram gets all flustered when I give it the boundary values?

OpenStudy (anonymous):

@Loser66 Wait are you saying \[ \frac{du}{dr}=\frac{d^2r}{(dt)^2} \]Or\[ \frac{du}{dt}=\frac{d^2r}{(dt)^2} \]

OpenStudy (anonymous):

Pick one.

OpenStudy (loser66):

last one

OpenStudy (anonymous):

That would mean: \[ du=u'dt \neq u'dr \]

OpenStudy (loser66):

oh yea, you are right,

OpenStudy (loser66):

surrender, It's toooooough.

OpenStudy (inkyvoyd):

uhm, this might not be the best time to mention this, but it just so happens I meant BVP and not Diff eq in my question

OpenStudy (anonymous):

Hmm, It makes me want to guess a trig function like \(\sin(t)\)

OpenStudy (anonymous):

But that doesn't really make sense does it.

OpenStudy (inkyvoyd):

I GOT IT.

OpenStudy (inkyvoyd):

the numerical solution I mean

OpenStudy (anonymous):

How?

OpenStudy (inkyvoyd):

wasn't using syntax correctly. But the solution is as expected (the problem I gave was what were to happen if one stopped the earth's orbit and watched as it hit the sun)

OpenStudy (anonymous):

Yeah it is the position of anything heading towards another mass with gravity

OpenStudy (anonymous):

It's not hard to see it sort of is negative exponential, but

OpenStudy (inkyvoyd):

It's plotting on mathematica for me, but I don't know whether or not it's not converging the way i want it to or I'm just bad